OpenStudy (anonymous):

hookes equation indicated F=kx where x is the stretch of a spring from equilibrium and k is the stiffness of the spring measured in kg/s^2. Given this, what is the SI unit of F?

OpenStudy (anonymous):

F=kx F=(kg/s^2 )(m) F=Kgm/s^2=N (Newton)